OpenStudy (anonymous):

slope intercept? i need help really fast. - - - FIND THE SLOPE AND Y-INTERCEPT OF THE GRAPH OF EACH EQUATION. - 1. y=3x-5 2. y= -5x +13 3. y= -x - 1 4. y= -11+ 6 5.y=-5 thanks

OpenStudy (anonymous):

ok so slope intercept is in the form of mx+b=y

OpenStudy (anonymous):

so the slope is mx and the y intercept is b

OpenStudy (anonymous):

yeah i know i just don't know how to do it. explain

OpenStudy (anonymous):

so for number 1 3 is the slope because it is m m is the slope

OpenStudy (anonymous):

so is number 2 -5 slope

OpenStudy (anonymous):

-5 is b and the y intercept -5 is tells u where the graph intersects with the y axis it intercepts at 0,-5

OpenStudy (anonymous):

can it be negative?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

technically it is 5/1 because all numbers are fractions

OpenStudy (anonymous):

but thats the same as a slope of 5

OpenStudy (anonymous):

are u getting it now?

OpenStudy (anonymous):

oh kinda

OpenStudy (anonymous):

for the last problem it says y=-5 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

this is a slightly different type of problem then the rest

OpenStudy (anonymous):

need to find m right

OpenStudy (anonymous):

ok so problem 3 you have -x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the slope is -1 because there is always a one in front of one variable

OpenStudy (anonymous):

for example x can also be written as 1x there fore when x is written as -x it means -1x

OpenStudy (anonymous):

because 1x is basically it self

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yep you got it!

OpenStudy (anonymous):

the last two problems don't have a slope they are horozontal lines

OpenStudy (anonymous):

4 and 5?

OpenStudy (anonymous):

yes because there is no x in the equation

OpenStudy (anonymous):

because there is only y=a number it means that the line stays at that y value/ height forever

OpenStudy (anonymous):

oops 4. y= -11x+ 6

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so in this one it says -11x + 6 so -11x is mx and 6 is b

OpenStudy (anonymous):

so -11 is your slope

OpenStudy (anonymous):

are you starting to see how this works?

OpenStudy (anonymous):

the number in the equation next to x is the slope intercept right

OpenStudy (anonymous):

it is the Y intercept

OpenStudy (anonymous):

the number u see added to the x is always the y intercept in linear equations

OpenStudy (anonymous):

*added or subtracted

OpenStudy (anonymous):

i thought that was he y- intercept y=mx+b <<<---- b

OpenStudy (anonymous):

the**

OpenStudy (anonymous):

yes B is the y intercept

OpenStudy (anonymous):

sorry if i confused u

OpenStudy (anonymous):

okay okay i got a little confused there

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

ok so are you cool with all of this or do u have more questions

OpenStudy (anonymous):

i mean i can try to finish it i kinda get it but i'm not like a pro yet, ya know

OpenStudy (anonymous):

but thanks tho

OpenStudy (anonymous):

dont worry practice at this stuff will make u better its always confusing at first

OpenStudy (anonymous):

wait i do have this problem farther in the chapter you can clearify

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the given point. 19. (3,5) and (0,4) 20. (2,6) and (-4,_2)

OpenStudy (anonymous):

points**

OpenStudy (anonymous):

ok so do u know how to calculate slope?

OpenStudy (anonymous):

Select one to be (x1, y1) and the other to be (x2, y2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok so this is the formula u need to know its rise over run so y1-y2/x1-x2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

when u use this formula it will give u your slope or m

OpenStudy (anonymous):

okay good so far

OpenStudy (anonymous):

After you calculate your slope you have your slope and your points So to find the y intercept you need to solve for b by setting up your equation like this First choose one of the points given to use in the equation to solve for b then put in the y coordinate in for the y in the equation then put your slope in the equation and solve for b

OpenStudy (anonymous):

ill demonstrate on the first one

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3,5 and 0,4 4-5/0-3=slope -1/-3=slope 1/3=slope y=mx+b y=5 x=3 m=1/3 b=? 5=1/3*3+b 5=1+b 4=b The Formula is y=1/3x+4

OpenStudy (anonymous):

oh my god.

OpenStudy (anonymous):

Ok you can see that I solved for the y intercept I didn't need to that for this problem because one of the coordinates given was the y intercept whenever you see a point with zero as the x value and a number for the y value remember that is the y intercept from there then u can easily get the b value

OpenStudy (anonymous):

u want me to explain more?

OpenStudy (anonymous):

do u understand the substitution

OpenStudy (anonymous):

I got the y and x values I used to find the B from one of the coordinates given to me

OpenStudy (anonymous):

okay on a scale of 1-10 how easy is this for you

OpenStudy (anonymous):

1 but i have been doing this a long time when i first did this it was hard

OpenStudy (anonymous):

dont worry things always start out tough

OpenStudy (anonymous):

is it like pypythagorean theorem easy? and 10 being easiest

OpenStudy (anonymous):

it depends if you have to square root any weird numbers

OpenStudy (anonymous):

mmm okay

OpenStudy (anonymous):

ohh 10 is easiest i though 1 was

OpenStudy (anonymous):

https://www.youtube.com/watch?v=u3spOO-m_Gg that video really helps but i feel like it has nothing to do with my assignment

OpenStudy (anonymous):

ill take a look

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok i didn't watch the whole thing (sorry i still have some of my own hw so I can't take all day) but I would highly recommend khan academy videos I know those are a little fast paced but they are really good ur video seemed ok it wasn't super clear based on what i read from the subtitles and clarity is important when learning math

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

so how are feeling about slope intercept form on a scale of 1 to 10 10 being completely lost and 1 being a pro at it

OpenStudy (anonymous):

6

OpenStudy (anonymous):

ok do understand a little better than before hopefully?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thats GOOD

OpenStudy (anonymous):

I'm glad I could help

OpenStudy (anonymous):

thanks buddy

OpenStudy (anonymous):

I have to get back to my HW so good luck on yours ok

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

okay how old are u anyways

OpenStudy (anonymous):

umm im kinda strict about giving imfo away on the internet

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

yeah guess your rightt. but thanks tho

OpenStudy (anonymous):

yeah good luck on future hw tests quizzes midterms and finals

OpenStudy (anonymous):

Bye:)

OpenStudy (anonymous):

see yah

OpenStudy (anonymous):

bye